OpenStudy (anonymous):

x^(x+6)=x^(4x-3) I got x=3, but there is another answer: x=1 why and how do I reach it?

OpenStudy (sheraz12345):

it should be 3 i guess

jhonyy9 (jhonyy9):

so how do you think 1 on indifferent exponent how many will be ?

jhonyy9 (jhonyy9):

hope yo know that 1 on indifferent exponent will be 1 yes ?

OpenStudy (anonymous):

That's a good point jhonyy, perhaps my example wasn't good enough, here is another one: (x-3)^(x-3)=(x-3)^(4-x)

OpenStudy (anonymous):

Again I get 3.5, but there is also 4

jhonyy9 (jhonyy9):

but for the first exercise the 3 is right too than you just need make try calcule it and see that will add the same answer

jhonyy9 (jhonyy9):

for the second you see that the base so in this case (x-3) is equale in other sides so than just you need make the exponents equales ok ?

OpenStudy (anonymous):

i have to make them equal to 1?

jhonyy9 (jhonyy9):

why 1 ?

OpenStudy (anonymous):

(x-3), if x is 4, then the base is 1 and therefore 1 to any power is 1

OpenStudy (anonymous):

(x-3)^(x-3)=(x-3)^(4-x) => x=4 (4-3)^(4-3)=(4-3)^(4-4) 1^1=1^0 1=1

jhonyy9 (jhonyy9):

yes 4 is right

OpenStudy (anonymous):

I see

jhonyy9 (jhonyy9):

hope you understand it now sure

OpenStudy (anonymous):

Yeah thanks man

jhonyy9 (jhonyy9):

my pleasure